### v1

@@ Appendix D Lemmas and Intermediary Results
Suppose Assumptions C.3 and C.2 are satisfied, then uniformly in $\ell\in\mathcal{L}$,
@@
where functions $m_{d}$ and $w$ are defined in Equations A.1, A.2 and A.3 and
@@
where the second equality holds because $m_{1}(\ell,\theta)$ is continuously differentiable in $\theta$ under Assumption C.2-(2), and the third equality is due to Assumption C.3.
@@
By Assumption C.2-(ii), the class of functions $\{1(p\leq P(Z,\theta)\leq p+r_{p}):\theta\in\Theta,p\in[0,1],r_{p}\in[0,1]\}$ is a Vapnik-Chervonenkis (VC) class of function. Therefore, the class of functions $\{1\{y\leq Y\leq y+r_{y}\}\times 1(p\leq P(Z,\theta):\theta\in\Theta,p\in[0,1],r_{p}\in[0,1],r_{y}\in[0,1]\}$ is also VC class. Hence, the process $\hat{\mathbb{G}}_{m_{1}}$ is stochastically equicontinuous with respect to $(\theta,\ell)$. Note $\hat{\theta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\theta_{0}$, then there exist $\delta_{n}\downarrow 0$ such that with probability approaching one, $(\hat{\theta},\ell)\in B((\theta_{0},\ell),\delta_{n})$, where $B((\theta_{0},\ell),\delta_{n})$ is a ball in $\Theta\times\mathcal{L}$ centered at $(\theta_{0},\ell)$ with radius $\delta_{n}$. Therefore,
@@
where the last equality is by the stochastic equicontinuity of the process $\hat{\mathbb{G}}_{m_{1}}$. Combine both Equations D.5 and D.1, the result then follows. $\boldsymbol{\Box}$
@@
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, C.3 and C.2 are satisfied, then uniform in $\ell$,
@@
$\hat{\nu}_{1}(\ell)=\hat{m}_{1}(y,r_{y},p_{2},r_{p},\hat{\theta})\cdot\hat{w}(p_{1},r_{p},\hat{\theta})-\hat{m}_{1}(y,r_{y},p_{1},r_{p},\hat{\theta})\cdot\hat{w}(p_{2},r_{p},\hat{\theta})$
To save space, for generic $\ell$, we write $\hat{m}_{1}(\hat{\theta})\equiv\hat{m}_{1}(\ell,\hat{\theta})$ and $\hat{w}(\hat{\theta})\equiv\hat{w}(\ell,\hat{\theta})$. Similarly, $m_{1}(\theta_{0})\equiv m_{1}(\ell,\theta_{0})$ and $w(\theta_{0})\equiv w(\ell,\theta_{0})$. Then,
$\hat{m}_{1}(\hat{\theta})\hat{w}(\hat{\theta})-$ $m_{1}(\theta_{0})w(\theta_{0})=(\hat{m}_{1}(\hat{\theta})-m_{1}(\theta_{0})+m_{1}(\theta_{0}))(\hat{w}(\hat{\theta})-w(\theta_{0})+w(\theta_{0}))-m_{1}(\theta_{0})w(\theta_{0})$
@@
Finally, under Assumption C.2, each element of $\triangledown_{\theta}m_{1}(y,r_{y},p,r_{p},\theta_{0})$ is Lipschitz continuous in $y$, $r_{y}$, $p$, $r_{p}$ and it implies that $\{\partial m_{1}(y,r_{y},p,r_{p},\theta_{0})/\partial\theta_{j}:(y,r_{y},p,r_{p})\in[0,1]^{4}\}$ is a VC class of functions for each $j$. Similarly, each element of $\triangledown_{\theta}w(p,r_{p},\theta_{0})$ is Lipschitz continuous in $p$, $r_{p}$. It follows that $\{\phi_{m_{1}}(y,r_{y},p,r_{p},\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$, $\{\phi_{m_{0}}(y,r_{y},p,r_{p},\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$ and $\{\phi_{w}(p,r_{p},\theta_{0}):(p,r_{p})\in[0,1]^{2}\}$ are all VC classes of functions. weak convergence follows from the fact that $\{\phi_{\nu_{0}}(y,r_{y},p_{1},p_{2},,r_{p},\theta_{0}):(y,r_{y},p_{1},p_{2},r_{p})\in[0,1]^{5}\}$ and $\{\phi_{\nu_{0}}(y,r_{y},p_{1},p_{2},,r_{p},\theta_{0}):(y,r_{y},p_{1},p_{2},r_{p})\in[0,1]^{5}\}$ are both VC classes of functions. Therefore, we have
@@
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, C.3, C.4 and C.2 are satisfied, then uniform in $\ell$ over $\mathcal{L}$,
$\sqrt{n}(\hat{\nu}_{1}^{b}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}^{b})-\hat{\nu}_{1}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}))$
@@
$\sqrt{n}(\hat{\nu}_{0}^{b}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}^{b})-\hat{\nu}_{0}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}))$
@@
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, C.3, C.4 and C.2 are satisfied, then $\hat{\sigma}^{2}_{d}(\ell)$ defined in (3.8) satisfies that for $d=0,1$, $\sup_{\ell}|\hat{\sigma}^{2}_{d}(\ell)-{\sigma}^{2}_{d}(\ell)|=o_{p}(1)$.
@@
Conditioning on the sample, because $W_{i}^{b}$ are i.i.d. across $b$ and $i$, has expectation and variance equal to one, we know $E[(W_{i}^{b}-1)(W_{j}^{b}-1)|(Y,D,Z)]=0$ and $E[(W_{i}^{b}-1)^{2}|(Y,D,Z)]=1$. As $B\rightarrow\infty$, the right-hand side converges in probability (with respect to the distribution of $\{W^{b}\}_{b=1}^{B}$) to $\frac{1}{n}\sum_{i=1}^{n}\phi^{2}_{\nu_{d},i}(\ell,\theta_{0})+o_{p}(1)$, which in turn converges to $\sigma^{2}_{d}(\ell)$ uniformly over $\ell$ as $n\rightarrow\infty$. $\boldsymbol{\Box}$
Appendix E The influence function with covariate case
@@
$\equiv\mathop{\mathrm{argmax}}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}D_{i}\log P(Z_{i},X_{i},\theta)+(1-D_{i})\log(1-P(Z_{i},X_{i},\theta)).$ (E.1)
@@
As in Appendix D, we make the following assumptions.
Assumption E.1
@@
$\sqrt{n}(\hat{\theta}-\theta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s_{\theta_{0}}(D_{i},Z_{i},X_{i},\theta_{0})+o_{p}(1),$ (E.2)
$\sqrt{n}(\hat{\beta}_{1}-\beta_{1})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s_{\beta_{1}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{1})+o_{p}(1),$ (E.3)
$\sqrt{n}(\hat{\beta}_{0}-\beta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s_{\beta_{0}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{0})+o_{p}(1),$ (E.4)
where $s_{\theta_{0}}(\cdot)$, $s_{\beta_{1}}(\cdot)$ and $s_{\beta_{0}}(\cdot)$ are measurable, satisfying $E[s_{\theta_{0}}(D_{i},Z_{i},X_{i},\theta_{0})]=0$,$E[s_{\beta_{1}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{1})]=0$, $E[s_{\beta_{0}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{0})]=0$, $E[\sup_{\theta}\|s_{\theta_{0}}(D_{i},Z_{i},\theta)\|^{2+\delta}]<\infty$, $E[\sup_{\beta}\|s_{\beta_{1}}(D_{i},Y_{i},Z_{i},X_{i},\beta)\|^{2+\delta}]<\infty$, and $E[\sup_{\beta}\|s_{\beta_{0}}(D_{i},Y_{i},Z_{i},X_{i},\beta)\|^{2+\delta}]<\infty$ for some $\delta>0$.
Note that under similar conditions as in Section 4 of Hsu, Liao and Lin (2022, Econometric Reviews), (E.3) and (E.4) would hold. We define the following quantities for generic $(y,r_{y},p,r_{p},b,\theta)$:
$m_{1}(y,r_{y},p,r_{p},b,\theta)=\mathbb{E}[D1(y\leq Y-X^{\prime}b\leq y+r_{y})1(p\leq P(Z,X,\theta)\leq p+r_{p})],$
$m_{0}(y,r_{y},p,r_{p},b,\theta)=\mathbb{E}[(D-1)1(y\leq Y-X^{\prime}b\leq y+r_{y})1(p\leq P(Z,X,\theta)\leq p+r_{p})],$
$w(p,r_{p},\theta)=\mathbb{E}[1(p\leq P(Z,X,\theta)\leq p+r_{p})].$
@@
By Assumption E.1, all elements of $\triangledown_{\theta}m_{1}(y,r_{y},p,r_{p},\beta_{1},\theta_{0})$, $\triangledown_{\beta}m_{1}(y,r_{y},p,r_{p},\beta_{1},\theta_{0})$, $\triangledown_{\theta}m_{0}(y,r_{y},p,r_{p},\beta_{0},\theta_{0})$, and $\triangledown_{\beta}m_{0}(y,r_{y},p,r_{p},\beta_{0},\theta_{0})$, are Lipschitz continuous in $y$, $r_{y}$, $p$, $r_{p}$, and each element of $\triangledown_{\theta}w(p,r_{p},\theta_{0})$ is Lipschitz continuous in $p$, $r_{p}$. It follows that $\{\phi_{m_{1}}(y,r_{y},p,r_{p},\beta_{1}\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$, $\{\phi_{m_{0}}(y,r_{y},p,r_{p},\beta_{0},\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$ and $\{\phi_{w}(p,r_{p},\theta_{0}):(p,r_{p})\in[0,1]^{2}\}$ are all VC classes of functions. Finally, let
@@
$\hat{\nu}_{1}(y,r_{y},p_{1},p_{2},r_{p},\hat{\beta}_{1},\hat{\theta})=\hat{m}_{1}(y,r_{y},p_{2},r_{p},\hat{\beta}_{1},\hat{\theta})\cdot\hat{w}(p_{1},r_{p},\hat{\theta})-\hat{m}_{1}(y,r_{y},p_{1},r_{p},\hat{\beta}_{1},\hat{\theta})\cdot\hat{w}(p_{2},r_{p},\hat{\theta}),$
$\hat{\nu}_{0}(y,r_{y},p_{1},p_{2},r_{p},\hat{\beta}_{0},\hat{\theta})=\hat{m}_{0}(y,r_{y},p_{2},r_{p},\hat{\beta}_{0},\hat{\theta})\cdot\hat{w}(p_{1},r_{p},\hat{\theta})-\hat{m}_{0}(y,r_{y},p_{1},r_{p},\hat{\beta}_{0},\hat{\theta})\cdot\hat{w}(p_{2},r_{p},\hat{\theta}).$
Lemma E.1
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, E.2, 3.3 and E.1 are satisfied, then,
$\sqrt{n}(\hat{\nu}_{1}(y,r_{y},p_{1},p_{2},r_{p},\hat{\beta}_{1},\hat{\theta})-\nu_{1}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{1},\theta_{0}))=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{\nu_{1},i}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{1},\theta_{0})+o_{p}(1),$ (E.5)
$\sqrt{n}(\hat{\nu}_{0}(y,r_{y},p_{1},p_{2},r_{p},\hat{\beta}_{0},\hat{\theta})-\nu_{0}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{0},\theta_{0}))=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{\nu_{0},i}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{0},\theta_{0})+o_{p}(1),$ (E.6)
@@
The proofs are similar to those in Appendix D, so we omit the details.

### v2

@@ Appendix D Lemmas and Intermediary Results
Suppose Assumptions B.3 and B.2 are satisfied, then uniformly in $\ell\in\mathcal{L}$,
@@
where functions $m_{d}$ and $w$ are defined in Equations A.2, A.3 and A.4 and
@@
where the second equality holds because $m_{1}(\ell,\theta)$ is continuously differentiable in $\theta$ under Assumption B.2-(2), and the third equality is due to Assumption B.3.
@@
By Assumption B.2-(ii), the class of functions $\{1(p\leq P(Z,\theta)\leq p+r_{p}):\theta\in\Theta,p\in[0,1],r_{p}\in[0,1]\}$ is a Vapnik-Chervonenkis (VC) class of function. Therefore, the class of functions $\{1\{y\leq Y\leq y+r_{y}\}\times 1(p\leq P(Z,\theta):\theta\in\Theta,p\in[0,1],r_{p}\in[0,1],r_{y}\in[0,1]\}$ is also VC class. Hence, the process $\hat{\mathbb{G}}_{m_{1}}$ is stochastically equicontinuous with respect to $(\theta,\ell)$. Note $\hat{\theta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\theta_{0}$, then there exist $\delta_{n}\downarrow 0$ such that with probability approaching one, $(\hat{\theta},\ell)\in B((\theta_{0},\ell),\delta_{n})$, where $B((\theta_{0},\ell),\delta_{n})$ is a ball in $\Theta\times\mathcal{L}$ centered at $(\theta_{0},\ell)$ with radius $\delta_{n}$. Therefore,
@@
[…]
@@
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, B.3 and B.2 are satisfied, then uniform in $\ell$,
@@
[…]
@@
Finally, under Assumption B.2, each element of $\triangledown_{\theta}m_{1}(y,r_{y},p,r_{p},\theta_{0})$ is Lipschitz continuous in $y$, $r_{y}$, $p$, $r_{p}$ and it implies that $\{\partial m_{1}(y,r_{y},p,r_{p},\theta_{0})/\partial\theta_{j}:(y,r_{y},p,r_{p})\in[0,1]^{4}\}$ is a VC class of functions for each $j$. Similarly, each element of $\triangledown_{\theta}w(p,r_{p},\theta_{0})$ is Lipschitz continuous in $p$, $r_{p}$. It follows that $\{\phi_{m_{1}}(y,r_{y},p,r_{p},\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$, $\{\phi_{m_{0}}(y,r_{y},p,r_{p},\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$ and $\{\phi_{w}(p,r_{p},\theta_{0}):(p,r_{p})\in[0,1]^{2}\}$ are all VC classes of functions. weak convergence follows from the fact that $\{\phi_{\nu_{0}}(y,r_{y},p_{1},p_{2},,r_{p},\theta_{0}):(y,r_{y},p_{1},p_{2},r_{p})\in[0,1]^{5}\}$ and $\{\phi_{\nu_{0}}(y,r_{y},p_{1},p_{2},,r_{p},\theta_{0}):(y,r_{y},p_{1},p_{2},r_{p})\in[0,1]^{5}\}$ are both VC classes of functions. Therefore, we have
@@
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, B.3, B.5 and B.2 are satisfied, then uniform in $\ell$ over $\mathcal{L}$,
$)\displaystyle\sqrt{n}(\hat{\nu}_{1}^{b}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}^{b})-\hat{\nu}_{1}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}$))
@@
$)\displaystyle\sqrt{n}(\hat{\nu}_{0}^{b}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}^{b})-\hat{\nu}_{0}(y,r_{y},p_{1},p_{2},r_{p},\hat{\theta}$))
@@
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, B.3, B.5 and B.2 are satisfied, then $\hat{\sigma}^{2}_{d}(\ell)$ defined in (3.6) satisfies that for $d=0,1$, $\sup_{\ell}|\hat{\sigma}^{2}_{d}(\ell)-{\sigma}^{2}_{d}(\ell)|=o_{p}(1)$.
@@
[…]
D.2 The influence function with covariate case
@@
$\equiv\mathop{\mathrm{argmax}}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}D_{i}\log P(Z_{i},X_{i},\theta)+(1-D_{i})\log(1-P(Z_{i},X_{i},\theta)).$ (D.11)
@@
As in Section D.1, we make the following assumptions.
Assumption D.1
@@
$\sqrt{n}(\hat{\theta}-\theta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s_{\theta_{0}}(D_{i},Z_{i},X_{i},\theta_{0})+o_{p}(1),$ (D.12)
$\sqrt{n}(\hat{\beta}_{1}-\beta_{1})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s_{\beta_{1}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{1})+o_{p}(1),$ (D.13)
$\sqrt{n}(\hat{\beta}_{0}-\beta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}s_{\beta_{0}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{0})+o_{p}(1),$ (D.14)
where $s_{\theta_{0}}(\cdot)$, $s_{\beta_{1}}(\cdot)$ and $s_{\beta_{0}}(\cdot)$ are measurable, satisfying $E[s_{\theta_{0}}(D_{i},Z_{i},X_{i},\theta_{0})]=0$$E[s_{\beta_{1}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{1})]=0$, $E[s_{\beta_{0}}(D_{i},Y_{i},Z_{i},X_{i},\beta_{0})]=0$, $E[\sup_{\theta}\|s_{\theta_{0}}(D_{i},Z_{i},\theta)\|^{2+\delta}]<\infty$, $E[\sup_{\beta}\|s_{\beta_{1}}(D_{i},Y_{i},Z_{i},X_{i},\beta)\|^{2+\delta}]<\infty$, and $E[\sup_{\beta}\|s_{\beta_{0}}(D_{i},Y_{i},Z_{i},X_{i},\beta)\|^{2+\delta}]<\infty$ for some $\delta>0$.
Note that under similar conditions as in Section 4 of Hsu, Liao and Lin (2022, Econometric Reviews), (D.13) and (D.14) would hold. We define the following quantities for generic $(y,r_{y},p,r_{p},b,\theta)$:
$m_{1}(y,r_{y},p,r_{p},b,\theta)=\mathbb{E}[D1(y\leq Y-X^{\prime}b\leq y+r_{y})1(p\leq P(Z,X,\theta)\leq p+r_{p})],$ (D.15)
$m_{0}(y,r_{y},p,r_{p},b,\theta)=\mathbb{E}[(D-1)1(y\leq Y-X^{\prime}b\leq y+r_{y})1(p\leq P(Z,X,\theta)\leq p+r_{p})],$ (D.16)
$w(p,r_{p},\theta)=\mathbb{E}[1(p\leq P(Z,X,\theta)\leq p+r_{p})].$ (D.17)
@@
By Assumption D.1, all elements of $\triangledown_{\theta}m_{1}(y,r_{y},p,r_{p},\beta_{1},\theta_{0})$, $\triangledown_{\beta}m_{1}(y,r_{y},p,r_{p},\beta_{1},\theta_{0})$, $\triangledown_{\theta}m_{0}(y,r_{y},p,r_{p},\beta_{0},\theta_{0})$, and $\triangledown_{\beta}m_{0}(y,r_{y},p,r_{p},\beta_{0},\theta_{0})$, are Lipschitz continuous in $y$, $r_{y}$, $p$, $r_{p}$, and each element of $\triangledown_{\theta}w(p,r_{p},\theta_{0})$ is Lipschitz continuous in $p$, $r_{p}$. It follows that $\{\phi_{m_{1}}(y,r_{y},p,r_{p},\beta_{1}\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$, $\{\phi_{m_{0}}(y,r_{y},p,r_{p},\beta_{0},\theta_{0}):(y,r_{y},p,r_{p})\in[0,1]^{4}\}$ and $\{\phi_{w}(p,r_{p},\theta_{0}):(p,r_{p})\in[0,1]^{2}\}$ are all VC classes of functions. Finally, let
@@
[…]
Lemma D.5
Suppose Assumptions 2.1, 2.4, 2.3, 2.2, D.2, 3.3 and D.1 are satisfied, then,
$\sqrt{n}(\hat{\nu}_{1}(y,r_{y},p_{1},p_{2},r_{p},\hat{\beta}_{1},\hat{\theta})-\nu_{1}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{1},\theta_{0}))=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{\nu_{1},i}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{1},\theta_{0})+o_{p}(1),$ (D.18)
$\sqrt{n}(\hat{\nu}_{0}(y,r_{y},p_{1},p_{2},r_{p},\hat{\beta}_{0},\hat{\theta})-\nu_{0}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{0},\theta_{0}))=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi_{\nu_{0},i}(y,r_{y},p_{1},p_{2},r_{p},{\beta}_{0},\theta_{0})+o_{p}(1),$ (D.19)
@@
The proofs are similar to those in Section D.1, so we omit the details.
Appendix E Additional Empirical Results
Table 4: FLL Semi-parametric test, B-Spline
2 Knots 3 Knots 4 Knots 5 Knots
$p_{f}$ $p_{s}$ combined $p_{f}$ $p_{s}$ combined $p_{f}$ $p_{s}$ combined $p_{f}$ $p_{s}$ combined
All 0.13 1.00 0.25 0.06 1.00 0.11 0.02 1.00 0.04 na 1.00 1.00
Aggressive Assault 0.02 1.00 0.04 0.01 0.91 0.02 0.00 0.96 0.00 na 1.00 1.00
Robbery 0.13 0.73 0.25 0.06 0.99 0.11 0.02 0.44 0.04 na 0.45 0.90
Drug Sale 0.18 0.83 0.36 0.09 0.33 0.18 0.03 0.55 0.06 na 0.73 1.00
Drug Possession 0.45 0.82 0.89 0.31 1.00 0.61 0.14 0.99 0.27 na 0.98 1.00